OpenStudy (anonymous):

Ill give a medal to who ever answers. [9.01] Determine whether the graph of y = −3x2 + 2x − 8 opens up or down and whether it has a maximum or minimum point. Up; Minimum Up; Maximum Down; Minimum Down; Maximum

OpenStudy (anonymous):

it is a quadratic, so a parabola the leading coefficient is negative (it is \(-3\)) and therefore it opens DOWN

OpenStudy (anonymous):

since it opens down, it has a maximum, but no minimum

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you mind answering another one?

OpenStudy (isaiah.feynman):

Oh no I just missed a medal!!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no i don't mind, go ahead and post

OpenStudy (anonymous):

[9.01] Identify the vertex for the graph of y = −3x2 + 6x − 2. (1, 1) (−1, 1) (1, −5) (−1, −5)

OpenStudy (anonymous):

I think it is (-1, -5)

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{6}{2\times (-3)}=1\]

OpenStudy (anonymous):

ok then, i still think im right on the -5 though

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

no, it is not \(-5\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if you let \(x=1\) you get \(-3+6-2\)

OpenStudy (anonymous):

so 1

OpenStudy (anonymous):

so vertex is \((1,1)\)

OpenStudy (anonymous):

ok thanks! :)

OpenStudy (anonymous):

yw